OpenStudy (anonymous):

Turn X+3y=12 And X=y-8 Into the formula y=mx+b

OpenStudy (anonymous):

Format the equations to solve for y. All you really need to do is add or subtract x then divid anything that is next to y onto the other side.

OpenStudy (anonymous):

Just reorder the values. x + 3y = 12 Subtract x from both sides. x +3y - x = 12 - x 3y = 12 - x Divide both sides by 3 so that y is completely isolated. 3y/3 = 12 - x /3 y = 4 - 1/3x Now move the slope so it's in slope intercept form. y= -1/3x + 4 x = y -8 Just add a positive 8 to both sides so y is isolated. x + 8 = y - 8 + 8 x + 8 = y Just turn into slope int. form again.